hi
i have gotten the answer to the question, but don't know the reasoning behind it. our teachers manual dosen't have the equation written out.
the question is as follows.
Find an angel such that 3 times the complement of the angel is 10 degrees less than the supplement of the angel.
please help

Answer by stanbon(75887)

Let the angle have measure "x".
Supplement is 180-x
Complement is 90-x
EQUATION:
3(90-x)=(180-x)-10
270-3x=170-x
2x=100
x=50 degrees